OpenStudy (anonymous):

Can anyone help me reduce 6/72?

OpenStudy (anonymous):

1/12 I believe is the answer. Just looked it up

OpenStudy (anonymous):

Hope that helps

jimthompson5910 (jim_thompson5910):

the GCF of 6 and 72 is 6. This is the largest factor in both numbers divide each term by the GCF 6/6 = 1 72/6 = 12 therefore, 6/72 = 1/12